OpenStudy (anonymous):

Simplify. x/6 1/6 fractions plz help

OpenStudy (anonymous):



OpenStudy (therealmeeeee):

ok 1/36 - 1/x^2 = (x^2 - 36) / 36x^2 = (x + 6)(x - 6) / 36x^2 1/6 + 1/x = (x + 6) / 6x So, [1/36 - 1/x^2] / [1/6 + 1/x] = [(x + 6)(x - 6) / 36x^2] / [(x + 6) / 6x] = (x + 6)(x - 6) / 36x^2 * 6x / (x + 6) [cancel out the common factors: (x + 6), 6x] = (x - 6) / 6x Edit: You can also multiply top and bottom by 36x^2. On the numerator, (1/36 - 1/x^2) * 36x^2 = x^2 - 36 = (x - 6)(x + 6) On the denominator, (1/6 + 1/x) * 36x^2 = 6x^2 + 36x = 6x(x + 6) So overall, (x - 6)(x + 6) / 6x(x + 6) = (x - 6) / 6x

OpenStudy (therealmeeeee):

do this help u

OpenStudy (anonymous):

yes thx

OpenStudy (anonymous):

i have 1 more i need help with can u help me

OpenStudy (therealmeeeee):

sure

OpenStudy (anonymous):